### v1

# The NOvA Collaboration

Abstract
The flux of cosmic ray muons at the Earth’s surface exhibits seasonal variations due to changes in the temperature of the atmosphere affecting the production and decay of mesons in the upper atmosphere. Using data collected by the NOvA Near Detector during 2018–2022, we studied the seasonal pattern in the multiple-muon event rate. The data confirm an anticorrelation between the multiple-muon event rate and effective atmospheric temperature, consistent across all the years of data. Previous analyses from MINOS and NOvA saw a similar anticorrelation but did not include an explanation. We find that this anticorrelation is driven by altitude–geometry effects as the average muon production height changes with the season. This has been checked with a CORSIKA cosmic ray simulation package by varying atmospheric parameters, and provides an explanation to a longstanding discrepancy between the seasonal phases of single and multiple-muon events.
I Introduction
The flux of muons produced in cosmic ray air showers and detected underground exhibits a well-known seasonal variation. This effect arises due to the competition between the decay and interaction of secondary mesons ($\pi,K$) produced in the primary cosmic-ray interaction. In the upper atmosphere, where most muons originate, higher temperatures in summer cause the atmosphere to expand, leading to lower density and a greater probability for mesons to decay into muons before interacting. Conversely, in winter, the atmosphere contracts, resulting in higher density, which increases the likelihood of mesons interacting before they can decay. As a result, the muon flux reaches a maximum in summer and a minimum in winter. In the past, this seasonal effect has been observed in multiple experiments Barrett et al. (1952); Sherman (1954); Castagnoli and Dodero (1967); Fenton et al. (1961); Y. Andreyev et al. () (Baksan Collaboration); Ambrosio et al. (1997); Bouchta ; Bellini et al. (2012); M. Selvi et al. () (LVD Collaboration); Desiati et al. ; S. Osprey et al. (2009) (MINOS Collaboration); Adamson et al. (2010, 2014); T. Abrahao et al. (2017) (Double Chooz Collaboration); F. P. An et al. (2018) (Daya Bay Collaboration). For our case, since the NOvA near detector is underground, the majority of the cosmic muons produced in the atmosphere do not reach the detector as the energy of the primary cosmic rays falls exponentially ($\sim E^{-2.7}$). Only a muon with an energy $E_{\mu}>\sec{\theta_{zen}}\times$ 50 GeV can reach the detector, where $\theta_{zen}$ is the muon zenith angle.
Large neutrino detectors are capable of detecting multiple-muon events, where several nearly parallel muon tracks from a single cosmic ray air shower arrive at the detector. While the seasonal effect of single muons is well understood, the behavior of multiple-muon events presented an intriguing anomaly. Previous studies, such as those from MINOS Adamson et al. (2015) and NOvA Acero et al. (2019), reported a winter maximum in observed multiple-muon rates, in direct contrast to the summer peak seen for single muons. To investigate this discrepancy, we conducted a detailed analysis of multiple-muon events using the NOvA Near Detector (ND). We concentrate on the size of the muon component of air showers compared to the size of our ND. Our approach incorporates a cosmic muon simulation using COsmic Ray SImulations for KAscade (CORSIKA) Heck et al. , along with four years of NOvA ND data. This study provides new insights into the longstanding puzzle of multiple-muon seasonal variations and offers an explanation through the altitude–geometry effect, described in Section III. Further sections describe the more recent NOvA ND data, temperature data, the CORSIKA simulation, cosine fits to the data, and a discussion of other seasonal effects.
II Previous data from MINOS and NOvA
The MINOS experiment first reported an unexpected winter maximum in multiple-muon events underground, contrasting with the well-established summer peak observed for single muons. This effect was observed in both the MINOS ND Adamson et al. (2015), located at 225 meters water equivalent (mwe), and the MINOS Far Detector (FD) Adamson et al. (2015), located at 2100 mwe underground. In the MINOS FD, the seasonal behavior was found to depend on the observed spatial separation of muons in an event, with closely spaced muons exhibiting a winter maximum while widely separated muons followed the expected summer maximum behavior. The MINOS study proposed that secondary hadronic interactions in the atmosphere played an important role in this result, though a definitive explanation was not presented.
The NOvA ND Acero et al. (2019) ($E_{\mu}~>$ 50 GeV), located at the same depth as the MINOS ND but with a different detector design, confirmed the winter peak in multiple-muon events in a dataset spanning 2015–2017. The NOvA FD Acero et al. (2021) ($E_{\mu}~>$ 1.5 GeV), situated at the surface with minimal overburden, also observed a strong seasonal dependence in multiple-muon showers, with a 30% stronger winter maximum at higher multiplicities compared to the lowest multiplicities and a 20% stronger effect for showers arriving from near the horizon compared to the most vertical showers. The previous NOvA ND study Acero et al. (2019) found that the strength of the seasonal variation, which was measured in three different ways (Rayleigh Power, correlation coefficient, and cosine fit), did not clearly depend on zenith angle, muon separation, or muon angular separation, but did noticeably depend on muon multiplicity.
The MINOS study explored four possible explanations for this effect Adamson et al. (2015), including the altitude–geometry effect, but ultimately discarded it based on the difference between winter and summer in the separation among muons in the detector.
In this paper, we extend the analysis of Ref. Acero et al. (2019) with four years of NOvA ND data, explain the seasonal variation of multiple-muon events with the altitude–geometry effect, and check that explanation with a CORSIKA Heck et al. based simulation of cosmic rays in the NOvA ND.
III The Altitude–Geometry Effect
The altitude–geometry explanation of the multiple-muon seasonal effect is based on the fact that a multiple-muon shower that originates higher in the atmosphere spreads out further than one originating lower. Thus, more muons from the shower miss a finite-size underground detector in summer than in winter. Although there would be more multiple-muon events in the summer, fewer of them would be detected as multiple-muon events because they originate higher in the atmosphere. This is the main explanation for the observed seasonal pattern of multiple-muon events with the MINOS Adamson et al. (2015) and NOvA Acero et al. (2019, 2021) detectors.
Figure 1: This shows how the radial spread of muons increases with increasing temperature. In summer, the atmosphere expands, so the muon birth altitude ($h$ = DF) increases compared to that in winter ($h_{0}$ = DE), and hence the radial spread also increases in summer ($r$) compared to winter ($r_{0}$) (figure from Ref. Kokoulin et al. (2016)).
The total mass of the atmosphere remains approximately constant throughout the year, but seasonal temperature differences cause vertical expansion in summer and contraction in winter. Under the assumption of an isothermal atmosphere, the ideal gas law $PV=nRT$ implies that a $\pm$2% change in absolute temperature results in a corresponding $\pm$2% shift in altitude for a given pressure level. For a shower origin at fixed pressure, the increased altitude in summer leads to a broader lateral spread of muons at the detector. In contrast, during winter, the lower altitude results in a more concentrated footprint as shown in Figure 1 Kokoulin et al. (2016). As a result, a pair of muons that would be identified as a multiple-muon event in winter may be detected as a single-muon event in summer, simply due to their larger separation at the detector level.
This is precisely what we observed in the CORSIKA simulation in section VII. Moreover, calculations by members of the IceCube collaboration based on the altitude–geometry effect seemed to explain the MINOS and NOvA multiple-muon seasonal data Gaisser and Verpoest (2021a, b). In Figure 2, it can be observed that the birth altitude of muons from multiple muon events is higher in summer than in winter. The mean altitude increases in the summer compared to the winter. As a result, the separation among muons in multiple-muon events also increases in summer compared to winter as seen in Figure 3.
In contrast, the single muon rate remains unaffected. For every single muon that misses a finite-sized detector because it starts at a higher altitude, there is another single muon that now hits the detector. There is no detector geometry effect for single muons: a consequence of muons starting at a higher altitude in the summer affects the observed muon multiplicity in a finite detector, but does not change the muon flux.
Figure 2: The production altitude distribution of multiple-muon events from CORSIKA with muon threshold energy of 50 GeV. Three winter months (Dec–Feb) and three summer months (Jun–Aug) are compared. Vertical lines represent the mean muon birth altitude for summer (red dashed line) and winter (black solid line).
Figure 3: Muon separation distribution of multiple-muon events from CORSIKA with muon threshold energy of 50 GeV. Three winter months (Dec–Feb) and three summer months (Jun–Aug) are compared. Vertical lines represent the mean muon separation for summer (red dashed line) and winter (black solid line).
It is instructive to consider the typical transverse momentum of particles in hadronic showers, and how that impacts the separation distribution of muons reaching a detector. The opening angle between a primary pion and the original cosmic ray varies as $\sim(p_{T}/p_{l})$, where $p_{T}$ is the transverse momentum of the pion and $p_{l}\sim E_{\pi}$ is the longitudinal momentum, relative to the direction of the primary cosmic ray. A typical separation for muons from two pions in the primary interaction would be $\sqrt{2}\times A\times(p_{T}/E_{min})$, where $A$ is the altitude where the pions decay and $E_{min}$ is larger for deeper detectors due to muon energy loss in the overburden. For a typical hadronic transverse momentum of $300\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ for two $60\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$ pions decaying at an altitude of $20\text{\,}\mathrm{k}\mathrm{m}$, one would expect a typical muon separation of $140\text{\,}\mathrm{m}$ at the NOvA ND. This is much larger than the size of the NOvA ND or MINOS ND or MINOS FD, so the altitude–geometry effect will decrease the number of observed multiple-muon events in the summer. In contrast, if the typical muon separation is smaller than the detector, which can be the case for deep detectors with TeV muon thresholds, the altitude–geometry effect is minimal. For example, the surface area of the MACRO experiment Ambrosio et al. (2002) is $76\text{\,}\mathrm{m}$$\times$$12\text{\,}\mathrm{m}$, but the threshold energy for an atmospheric muon to reach the detector is $1.3\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$. At this energy, the average separation among muons is approximately $8\text{\,}\mathrm{m}$, which is smaller than the size of the MACRO detector. Hence, MACRO observed a summer maximum for multiple-muon events Ronga (2017), just as for single muon events. However, in the MINOS FD, the typical muon separation is comparable to the size of the detector. This helps to explain why the observed muon separation in the detector determines whether there is a summer maximum or a winter maximum.
Every detector has a unique size, but there is a straightforward relationship between the depth of an underground detector and the minimum muon energy required to reach that detector. A detector at or near the surface can detect muons of a few GeV, and the size of an extended air shower can exceed a kilometer. At the MINOS FD, the muon energy threshold is near $0.7\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$, and the typical muon separation is a few meters. In other words, as detector depth increases, the muons that survive are coming from a smaller and smaller area of the core of the shower.
The role of summer/winter differences in the origin altitude of muons in an air shower, and its effect on shower containment in a finite detector, has been considered before by MINOS Adamson et al. (2015). However, the MINOS paper’s conclusion, discarding the effect, was based on the measured muon separation distributions in the near and far detectors without taking into account that the measured distribution does not reflect the true separation if the size of the muon shower is larger than the detector. The altitude–geometry effect was considered responsible for the seasonal results of the DECOR experiment Yurina et al. (2019), which measured muons close to the surface of the earth. For those muons, with an energy typically of a few GeV, other effects such as pressure changes are important, as well as muon decay, which is discussed further in Section VIII.
IV COSMIC MUON DATA IN THE NOvA NEAR DETECTOR
[…]
In each cell, the light emitted by the scintillator is collected and transported to an avalanche photodiode (APD) through a wavelength shifting fiber. The APD converts the light to an electronic signal, digitized by the front-end electronics, and recorded as a hit if the signal exceeds a preset threshold value.
Cosmic muons in the NOvA ND are collected using a data-driven activity trigger. This trigger initiates data collection when there are at least 10 hits in detector planes, requiring hits on at least 3 planes in each of the two views and a minimum of 8 planes in total. Additionally, hits must be recorded in at least 5 planes out of 6 consecutive planes.
Collected hits are reconstructed using the Hough Transform Ballard (1981) method which is widely used for line reconstruction. All possible reconstructed tracks in two dimensions (2D) are stored. These 2D tracks from each plane are merged to create a 3D track.
In order to ensure a clean sample of cosmic ray induced multiple muon events, three selection cuts are applied to the tracks. First, the track’s start and endpoint should be less than 50 cm from the detector surface. This fiducial cut eliminates a number of events in which an electron emerged from the rock above the cavern, nearly parallel to an accompanying muon, and made a short track. This cut also eliminates contained and partially contained muon events. Each muon track is required to cross at least 10 planes. Some single muon events that occur closely spaced in time are recorded as multiple-muon events. Hence, the time between the first and any other tracks of a multiple-muon event is limited to a maximum of $100\text{\,}\mathrm{n}\mathrm{s}$. This eliminates those falsely recorded as multiple-muon events. The fiducial cut removes 16.9% of the total number of events considered in the analysis. The selection criteria on the number of planes further eliminate 16.2% of the total events, and finally the timing cut removes an additional 0.04% of the events.
The track multiplicity distribution of selected events from the ND data is shown in Figure 4. The maximum multiplicity observed in the ND data is 10. The distribution of cosine of the zenith angle, representing the angle between the muon track and the vertical upward direction, is illustrated in Figure 5. The distribution exhibits good agreement between data and Monte Carlo simulation.
Figure 4: Distribution of number of reconstructed muon tracks in an event (multiplicity) in the dataset (2018–2022) used for this analysis. The highest multiplicity observed is 10. Error bars represent statistical uncertainties.
Figure 5: The distribution of the cosine of the angle between the track and the vertical upward direction (zenith angle). The ratio of Data over the CORSIKA MC prediction is shown in the lower panel with statistical uncertainty.
The multiple-muon event rate in the data is obtained by fitting the distribution of time difference ($\Delta t$) between pairs of consecutive events. It eliminates the effects of any data acquisition glitches, such as brief pauses resulting in data loss, or bursts of noise. The time differences for such cases appear in the outliers of the distribution while the core of the distribution is unaffected. The distribution is fitted with the Erlang distribution function,
$R[\lambda,x\equiv\log_{10}(\Delta t)]=Ae^{-\lambda 10^{x}}10^{x}\lambda.$ (1)
Here, $\Delta t$ is the time difference between consecutive events, $\lambda$ is the event rate, and $A$ is a normalization constant. The fitted distribution is shown in Figure 6 in the range 1.5 to 4.7, and the $\chi^{2}$/NDF from the fits as a function of time are shown in Figure 7. This is an improvement over the previous method used in the NOvA ND analysis Acero et al. (2019) that calculated the rate by dividing the total number of muons by the total exposure time.
Figure 6: Logarithm (base 10) of time difference between $i$ and $(i+1)$th multiple-muon events from one day of data. The distribution is fitted with an Erlang distribution in the range 1.5 to 4.7. Error bars represent statistical uncertainties.
Figure 7: $\chi^{2}$/NDF distribution of Erlang fit for multiple-muon data in weekly binning.
V Temperature data
Atmospheric temperature data are taken from the European Centre for Medium-Range Weather Forecasts (ECMWF) Dee et al. (2011). For this analysis, we utilize temperature data from ERA5 Hersbach et al. (2020), the fifth-generation atmospheric reanalysis produced by ECMWF. They provide temperature readings at 37 pressure levels, distributed nonuniformly from 1 to 1000 hPa. For comparison with the NOvA ND data, temperatures are extracted from ECMWF within the grid area ranging from (41N, 87W) to (42.5N, 88.5W).
Since we do not know the exact altitude at which each muon is produced in the atmosphere, effective temperature is used instead of exact temperature. This is essentially a weighted average of the atmospheric temperature, where the weights reflect the probability distributions of meson (mainly pion and kaon) production and decay into muons at different altitudes. The effective temperature is defined as Grashorn et al. (2010)
$T_{e\!f\!f}=\frac{\sum_{i=1}^{n}{T(P_{i})[W_{\pi}(P_{i})+W_{K}(P_{i})]\Delta P_{i}}}{\sum_{i=1}^{n}[W_{\pi}(P_{i})+W_{K}(P_{i})]\Delta P_{i}}.$ (2)
Here, $T(P_{i})$ is the temperature at pressure level $i$ ($P_{i}$), and $W_{K}(P_{i})$ and $W_{\pi}(P_{i})$ are the weights responsible for the contribution as a function of pressure (altitude) for those $K$ and $\pi$ decays that make muons capable of reaching the NOvA ND. The effective temperature is calculated using cross sections and lifetime values from Grashorn et al. (2010), along with temperature values obtained at 37 pressure levels from ECMWF. Figure 8 shows the variation of ECMWF temperature data of the atmosphere over the Fermilab area, averaged over the four years 2018–2022, as a function of altitude and pressure. It also shows the total weight $W_{\pi}+W_{K}$ used in Equation 2.
Figure 8: Variation of average temperature with altitude and total weight function used to calculate effective temperature.
VI Cosmic muon simulation
Cosmic muons are simulated with CORSIKA Heck et al. v7.7410 using GHEISHA as a low-energy hadronic interaction model and QGSJET as a high-energy hadronic interaction model. The energy of primary particles is chosen to be between $200$ and $10^{6}$ GeV following a power law with exponent 2.7-2.7 Navas et al. (2024). Each muon in a multiple-muon event is required to have an energy above $50\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$, which is necessary to reach the NOvA ND vertically. A muon coming at a high zenith angle has to pass through a larger overburden and hence requires more energy to reach underground detector compared to vertically incident muons. This is taken care of by calculating the energy loss using Equation 3. Cosmic showers are simulated using the Earth’s magnetic field for the Fermilab location (41.84N, 88.25W).
The energy loss of muons due to the overburden above the ND is calculated using the relation
$E_{\mu}=(E_{\mu,0}+\epsilon)e^{-bd\rho}-\epsilon,$ (3)
where $E_{\mu,0}$ and $E_{\mu}$ denote the initial and final energies of the muons, respectively, as they traverse a distance $d$ through the underground rock with density $\rho$. The parameters $\epsilon$ and $b$ are material-dependent constants, with values of $500\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$ and $4\times 10^{-6}~\text{cm}^{2}/\text{g}$, respectively, for standard rock Zyla et al. (2020). Muons that retain nonzero energy after passing through the overburden are considered for the rate calculation below.
CORSIKA models the change in atmospheric density with altitude using a five-layer structure that extends from the ground to the top of the atmosphere ($100\text{\,}\mathrm{k}\mathrm{m}$). The density in the bottom four layers decreases exponentially with altitude, described in terms of the atmospheric mass overburden $X(h)$ and height $h$ by
$X(h)=a_{i}+b_{i}e^{-\frac{h}{c_{i}}},$ (4)
where $i$ ranges from 1 to 4. As the atmosphere becomes exceedingly thin in the top layer, the mass overburden decreases linearly with height for the 5th layer:
$X(h)=a_{5}-b_{5}h/c_{5}.$ (5)
CORSIKA provides default atmospheric profiles for some selected locations but does not support the Fermilab site. To address this, we extracted vertical atmospheric profile data during 2017 from the Global Data Assimilation System (GDAS) National Centers for Environmental Prediction (2024) (NCEP) and matched CORSIKA’s exponential layer model. The site-specific atmospheric coefficients ($a_{i}$, $b_{i}$, and $c_{i}$) are used as input parameters to the simulation. Proton showers are generated using simulation parameters described earlier in this section and atmospheric parameters extracted from GDAS. We found good agreement between the simulation and the data collected by the NOvA experiment in the ND, as shown in  Figure 5.
In the study of the seasonal effect, the single and multiple-muon event rates are calculated by using the events on a daily basis, and later binned by month. The passage of events through NOvA ND is simulated using GEANT Agostinelli et al. (2003); Allison et al. (2006, 2016).
After the generation of showers, cosmic muons from each generated shower are spread uniformly (randomized) over an area from $32\text{\,}\mathrm{m}$ to $32\text{\,}\mathrm{m}$ along the $x$-axis (horizontal) and $30\text{\,}\mathrm{m}$ to $46\text{\,}\mathrm{m}$ along the $z$-axis (along the neutrino beam) of the NOvA coordinate system. For multiple-muon events, the first muon is randomized uniformly over the $xz$ surface, and the remaining muons are placed on the $xz$ surface without changing the distance between muons and orientation given by CORSIKA. The ND’s surface area spans from $2\text{\,}\mathrm{m}$ to $2\text{\,}\mathrm{m}$ along the $x$-axis and from 0 to $16\text{\,}\mathrm{m}$ along the $z$-axis; the extended area for randomization is used to allow muons to hit the side of the detector. This choice allows us to include the core of 99.3% of all cosmic muons that hit the detector, and is much larger than the actual detector. Showers that produce at least one muon in the ND proceed to the detector simulation. Neutrino-induced upward-going muons are not considered for this analysis.
We studied the seasonal results for muons from two distinct stages of the simulation. We show results for all muons in a CORSIKA event that have sufficient energy to reach the underground detector. We refer to these results as predictions for an infinite detector, for which there was no detector simulation. We also used the CORSIKA output to simulate events in the NOvA ND as described in the previous paragraph. In this case, the seasonal results were obtained with the same software as used for NOvA ND data Acero et al. (2020). It was the marked difference in the results before and after the detector simulation that supported the hypothesis that a geometry effect was involved.
We do not claim that our implementation of CORSIKA accurately represents the altitude and temperature distributions in the summer and winter. In fact, the magnitude of the seasonal modulation in our CORSIKA simulation does not match the expected magnitude based on the well-established theoretical and experimental understanding of single-muon seasonal behavior Grashorn et al. (2010). However, we do expect the density profile in the atmosphere to provide a fairly good basis for the summer/winter differences regarding the hadronic showers and the competition between hadron interaction and decay in the upper parts of the atmosphere. Our use of the simulation in this analysis relies only on a comparison of the seasonal phases and magnitude for single- and multiple-muon events in a detector of fixed size.
VII Fits to Seasonal Variations
The simulated multiple-muon rate for an infinite detector shows a summer maximum as displayed in Figure 9, opposite that seen in multiple-muon measurements. When the simulated infinite detector events were passed through the finite ND, the time of the seasonal maximum moved from summer to winter, as shown in Figure 10. This result from our simulation is consistent with previous data observations (NOvA Acero et al. (2019) and MINOS Adamson et al. (2015)).
Figure 9: Seasonal variation of multiple-muon rate for an infinite detector at the depth of the NOvA ND from CORSIKA initiated by primary proton cosmic rays. The solid curve shows a cosine fit.
Figure 10: Seasonal variation of multiple-muon rate at the NOvA ND (finite detector) from CORSIKA initiated by primary proton cosmic rays. The solid curve represents a cosine fit of the variation. Error bars represent statistical uncertainties.
Figure 11 shows the multiple-muon event rate seen in the four-year (May 2018 to April 2022) NOvA analysis sample discussed above, as well as the effective temperature. Both the data and the temperature are presented in weekly bins. The percentage variation is calculated using
$\frac{\Delta R_{\mu}}{\langle R_{\mu}\rangle}=\frac{(R_{\mu}-\langle R_{\mu}\rangle)}{\langle R_{\mu}\rangle}$ (6)
where $R_{\mu}$ is the multiple-muon rate and $\langle R_{\mu}\rangle$ is the average rate over the four years of data. Systematic errors on the calculation of the rate, temperature and fit parameters are presented in Acero et al. (2019). None of these systematic errors varies as a function of season.
The muon rate in the multiple-muon data displays a seasonal trend that is out of phase with the effective atmospheric temperature. Although periodic, there is no a priori reason to expect a cosine shape. However, it does serve to extract the phase of the magnitude maxima, so the percentage variations of multiple-muon rate and effective temperature are fitted with cosine curves:
$V_{0}+V\cos\left[{\frac{2\pi}{t_{year}}(t-\phi)}\right],$ (7)
where $V_{0}$ is the average rate, $V$ is the magnitude, $t_{year}$ is the time period (fixed at one year) and $\phi$ is the phase. The fitted parameters are shown in Table 1. The multiple-muon data vary by $\pm 2\%$, and have the opposite phase as the temperature.
Table 1: Best fitted parameters for the cosine fit of percentage variation of the multiple-muon data and effective temperature
Muon data Temperature data
$V_{0}$ (%) -0.060 $\pm$ 0.002 0.050 $\pm$ 0.006
Magnitude ($V$) (%) 2.09 $\pm$ 0.04 1.11 $\pm$ 0.01
[…]
Figure 11: Seasonal variation in weekly binning: The percentage variation of observed multiple-muon rate (top) with cosine best fit (solid curve) and the percentage variation of effective temperature (bottom) with cosine best fit (solid curve).
VIII Other Seasonal Effects
In addition to the aforementioned altitude–geometry effect, the MINOS paper Adamson et al. (2015) discussed three more ideas to explain the observed seasonal pattern of multiple-muon event rate: (1) hadronic dimuon decays, (2) a temperature effect in layers of the atmosphere, and (3) an anticorrelation of primary and secondary meson decays. Another possible explanation is muon decay. We next discuss each in turn.
(1) Although the decay probability of pions increases in the summer, the decay probabilities of other hadrons, such as $\eta$ and $\rho$ mesons that decay into dimuons, change negligibly. However, due to the small dimuon branching ratio of these mesons (branching ratios for $\rho\rightarrow\mu^{+}\mu^{-}$ and $\eta\rightarrow\mu^{+}\mu^{-}\gamma$ are $4.6\times 10^{-5}$ and $3.1\times 10^{-4}$, respectively) hadronic dimuon decays are not sufficient to explain the opposite behavior in multiple-muon event rate compared to single muons.
(2) The temperature effect mentioned in Adamson et al. (2015) implies that there could be a seasonal variation dependent on altitude, differing for single and multiple-muon event rates. In the MINOS paper Adamson et al. (2015), the temperature data were examined to see any significant temperature changes, leading to changes in rate, at a particular altitude, but no such effect was observed.
(3) In the MINOS paper Adamson et al. (2015), the most plausible explanation for the observed phenomenon was considered to be the anticorrelation between the primary and secondary mesons. According to this hypothesis, the presence of multiple muons may arise not solely from the decay of mesons generated in the initial interaction, but also from mesons produced in subsequent interactions involving secondary mesons and air molecules. The higher probability of meson interactions in winter was posited as a contributing factor, leading to an increased number of mesons that, upon decay, yield muons and consequently contribute to the occurrence of more multiple-muon events. If this were true, the multiple-muon rate should exibit a winter maximum in the infinite detector. This explanation is contradicted by our simulation results; in Figure 9 we observed a summer maximum.
(4) Another potential explanation for the observed seasonal pattern of multiple muons is muon decay. Due to relativistic effects, this phenomenon is most pronounced at low muon energies, as observed by detectors at or near the surface. The seasonal effect of muon decay for various muon thresholds can be found in Table 2. Surface detectors such as the NOvA far detector Acero et al. (2021), DECOR Yurina et al. (2019); Tolkacheva et al. (2011), and GRAPES Arunbabu et al. (2017), all with muon thresholds near 5 GeV, can exhibit a 0.6% excess of events during the winter. In contrast, detectors such as the NOvA ND and MINOS ND, which have muon thresholds of 50 GeV, would only show a smaller 0.09% excess in winter. This effect is even smaller for the MINOS FD and MACRO.
Table 2: A prediction of the percentage of muons that reach sea level before they decay as a function of muon energy. $E_{\mu}$ is the muon energy, $N_{0}$ is the initial number of muons, and $N$ is the number of muons that survive. The summer/winter altitude difference is assumed to be 2%.
$E_{\mu}$ (GeV) $N/N_{0}$ (%) $N/N_{0}$ (%) Difference (%)
(winter) (summer)
5 62.05 61.46 0.59
50 95.34 95.25 0.09
500 99.52 99.51 0.01
IX Conclusion
The NOvA ND multiple-muon event rates have been measured using the Erlang fitting technique, which eliminates systematic effects due to detector and DAQ instabilities. The measured multiple muon rate for the 2018 to 2022 data period is found to be anticorrelated with the effective temperature of the atmosphere. This anticorrelation is consistent with previous NOvA ND and MINOS measurements.
The altitude–geometry effect is the most relevant explanation for the opposite seasonal behaviors between single and multiple-muon events. Evidence supporting this conclusion comes from a comparison between the typical muon separation distribution as a function of underground detector depth vis-a-vis the detector size, as well as a CORSIKA simulation, which demonstrates a summer maximum for multiple-muon events in an infinite detector. However, for a finite-sized detector, such as the NOvA ND, the simulation shows a winter maximum for multiple-muon events. Other proposed solutions would have resulted in a winter maximum in the CORSIKA simulation for an infinite detector. Additionally, the explanation of muon decay was found to be too small to account for observations of $50\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$ muons. The altitude–geometry effect was found to be consistent with observations from the MINOS FD, MINOS ND, NOvA ND, MACRO, and NOvA FD. All other considered explanations were excluded on multiple grounds.
Acknowledgements.
This document was prepared by the NOvA collaboration using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, HEP User Facility. Fermilab is managed by Fermi Forward Discovery Group, LLC, acting under Contract No. 89243024CSC000002. This work was supported by the U.S. Department of Energy; the U.S. National Science Foundation; the Department of Science and Technology, India; the European Research Council; the MSMT CR, GA UK, Czech Republic; the RAS, the Ministry of Science and Higher Education, and RFBR, Russia; CNPq and FAPEG, Brazil; UKRI, STFC and the Royal Society, United Kingdom; and the State and University of Minnesota. We are grateful for the contributions of the staffs of the University of Minnesota at the Ash River Laboratory, and of Fermilab. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.